OpenStudy (anonymous):

If the graph of an equation is symmetric with respect to the origin and (3,-4) is a point on the graph then what is also a point on the graph?

OpenStudy (debbieg):

Do you know what it means for a function to be symmetric wrt to the origin? What does that mean, algebraically?

OpenStudy (anonymous):

^exactly what does it mean?

OpenStudy (debbieg):

If a function f(x) is symm wrt the origin, it means that: \[\large f(-x)=-f(x)\]for all x in the domain of f(x)

OpenStudy (debbieg):

An example would be \(\large f(x)=x^3\), since for any x, if I compare the y value at x and the y value at -x, they are opposites. E.g. \(\large f(2)=8\) and \(\large f(-2)=-8\) So can you take that definition, and relate it to what it means regarding the point given above?

OpenStudy (anonymous):

What about the y coordinate then?

OpenStudy (debbieg):

Huh? y=f(x) f(x) IS the y-coordinate. Every point on the graph is (x, f(x))

OpenStudy (anonymous):

Good luck

OpenStudy (debbieg):

In the cubic example I gave above, the 2 points are (2,8) and (-2,-8)

OpenStudy (anonymous):

So then for it to be symmetric with respect to the origin would the point be (-3,4)?

OpenStudy (anonymous):

one example as my teacher gave which i liked is the ying yang sign - can be rotated 180 degrees is the symmetric with respect to the origin

OpenStudy (anonymous):

 do you see?

OpenStudy (anonymous):

yes I do! that helped so much I thought that's what it was! thank you so much!

OpenStudy (anonymous):

:) taking precalculus now so glad i could help :)

OpenStudy (anonymous):

it was a ton of help! good luck with that :))

OpenStudy (debbieg):

Yes, (-3,4), you got it! :)

OpenStudy (anonymous):

yay! :)

OpenStudy (anonymous):

does anyone know the intercepts to y=5x^2-5

OpenStudy (anonymous):

to find the x and y intercepts set y=0 then solve for x then set x to 0 and solve for y

OpenStudy (anonymous):

I got that one too! thank you again! :)